OpenStudy (toxicsugar22):

Duncan knows that it takes 36400cal of energy to heat a pint of water from room temperature to boiling. However, Duncan has prepared ramen noodles so many times he does not need to measure the water carefully. If he happens to heat 0.600pint of room-temperature water, how many kilojoules of heat energy will have been absorbed by the water at the moment it begins to boil?

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

So It takes 1 pint of water 36,400 cal of energy to go from room temperature to boiling

OpenStudy (anonymous):

If he heats 0.6 pints of water, how many kj of heat energy were absorbed? Do you know how to convert calories to joules?

OpenStudy (anonymous):

Did you guys ever use 4.184 Joules = 1 calories?

OpenStudy (anonymous):

Nope, got to do homework myself. I can't do all the work, sorry. It's really easy though, you can figure it out

OpenStudy (toxicsugar22):

4.184 Joules = 1 calories?

OpenStudy (toxicsugar22):

cdo you understand this

OpenStudy (theeric):

I found that 4.186 joules make one calorie, so I'm guessing it is around there. Or 238.9 calories for one kilojoule. @toxicsugar22 What I think you need to do, but I'm not completely sure (I haven't studied chemistry in a while), is that you take 36,400 calories to heat up one pint. For .6 pint, you only need .6/1 of the calories. Then, you will have the number of calories used to boil the water, and you'll need to convert that to kilojoules. Have you ever worked with making a conversion factor to convert units?

OpenStudy (theeric):

To get a conversion factor, you find two equal amounts, like \(238.9\ [cal]= 1\ [kJ]\). Then you divide both side by one side, like \(\dfrac{238.9\ [cal]}{238.9\ [cal]}=1= \dfrac{1\ [kJ]}{238.9\ [cal]}\) Now \(1= \dfrac{1\ [kJ]}{238.9\ [cal]}\), and you can multiply anything by \(1\) without changing the actual amount! Which is GREAT for converting. You can multiply it by a number of calories and the calorie units will drop and leave only kilojoules. And the actual amount has not changed, since you multiplied it by \(1\).

OpenStudy (toxicsugar22):

not really this is my first time

OpenStudy (theeric):

First time doing what?

OpenStudy (toxicsugar22):

this problem

OpenStudy (toxicsugar22):

can you explain to me ever procedure

OpenStudy (theeric):

Okay. So, read what I posted before, and tell me what part first confused you: "What I think you need to do, but I'm not completely sure (I haven't studied chemistry in a while), is that you take 36,400 calories to heat up one pint. For .6 pint, you only need .6/1 of the calories."

OpenStudy (toxicsugar22):

ok i got that part

OpenStudy (toxicsugar22):

now what next

OpenStudy (toxicsugar22):

ok clear on everthing

OpenStudy (theeric):

You got it? :)

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

so tell me the rest now

OpenStudy (theeric):

So you know how many calories it takes to heat up .6 pint. Now, you convert that to kilojoules. You can use a "conversion factor," which is just a number you multiply with to convert units. I made one up above! Now read that post, and let me know if you have a question!

OpenStudy (toxicsugar22):

1 calorie it takes for .6 pint

OpenStudy (toxicsugar22):

is that right

OpenStudy (toxicsugar22):

im taking notes from you

OpenStudy (theeric):

I think that you'll have .6 of the amount of water, you'll need .6 the amount of calories! So you'll multiply the number of calories used to heat 1 pint of water by .6. My thinking is that the energy used to boil water is always a proportion to the amount of water. \(\dfrac{136,400\ calories}{1\ pint} = some\ ratio\) And the ration of calories to amount of water is the same, I think. So \(\dfrac{136,400\ calories}{1\ pint} =\dfrac{x\ calories}{.6\ pint}\) Solve for "\(x\ calories\)," because that is how many calories it takes to heat the water! I have to go! Good luck!

OpenStudy (toxicsugar22):

would it be 227333.33

OpenStudy (toxicsugar22):

Lethal would this be it

OpenStudy (theeric):

You would multiply both sides by \(.6\ pint\). It seems like you divided. And then you have the number of calories and you need, and you just need that in kilojoules. You would multiply it by a conversion factor, then. I wrote one up above, when I talked about the conversion factor.

OpenStudy (anonymous):

It's too easy

OpenStudy (anonymous):

You can't miss it